1. Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the NPV of the PJX5?

a. The PJX5 will cost $1.63 million fully installed and has a 10 year life. It will be depreciated to a book value of $241,344.00 and sold for that amount in year 10.

b. The Engineering Department spent $49,388.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $15,137.00.

d. The PJX5 will reduce operating costs by $469,222.00 per year.

e. CSD’s marginal tax rate is 33.00%.

f. CSD is 57.00% equity-financed.

g. CSD’s 12.00-year, semi-annual pay, 6.47% coupon bond sells for $972.00.

h. CSD’s stock currently has a market value of $24.59 and Mr. Bensen believes the market estimates that dividends will grow at 4.44% forever. Next year’s dividend is projected to be $1.42.

2. Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the NPV of the PJX5?

a. The PJX5 will cost $1.63 million fully installed and has a 10 year life. It will be depreciated to a book value of $272,465.00 and sold for that amount in year 10.

b. The Engineering Department spent $15,762.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $18,418.00.

d. The PJX5 will reduce operating costs by $316,152.00 per year.

e. CSD’s marginal tax rate is 25.00%.

f. CSD is 55.00% equity-financed.

g. CSD’s 13.00-year, semi-annual pay, 6.20% coupon bond sells for $1,033.00.

h. CSD’s stock currently has a market value of $21.02 and Mr. Bensen believes the market estimates that dividends will grow at 3.43% forever. Next year’s dividend is projected to be $1.55.

Answer format: Currency: Round to: 2 decimal places.
Answer #1

1.

Depreciation = Cost - Salvage Value /no. of life

Depreciation=

(1630000-241344)/10

Depreciation=

              138,866

Now we need to find the WACC from the information given to use as required rate of return

WACC= {kd (1-t)*debt/ debt+ equity}+ {ke*equity/debt+ equity}

So we need to find the cost of equity and cost of debt:

cost of old Equity

= Dividend for next year/ equity share price+ growth

growth

4.40%

dividend of next yr

1.42

Cost of equity=

=1.42/24.59 +4.4%

0.057+0.044

0.1017

=10.17%

Cost of Bond

YTM= (C+ (F-P)/n)/(F+P/2)

C= coupon amount= 1000*6.47= 64.7

F= face value=1000

P= Price= 972

N= tenor= 12

YTM= (64.7+(1000-972)/12)/(1000+972/2)

YTM=67.033/986

YTM= 0.6798 or 6.798%

Particulars

Cost

tax

After tax cost

weight

after tax cost * weights

Bonds

6.80%

34%

=0.068*(1-0.34 )

= 0.04488 or 4.49%

43%

0.0449*0.43

0.0193

Equity

10.17%

0%

10.17%

57%

0.1017*0.57

     0.06

total

0.0773

Weighted average cost of capital = 7.73%

Now using this as required rate of return we will find the NPV of the project.

So ,

Year

Particulars ( cash flows)

cash flow

Depreciation

profit before tax ( income-dep.)

Tax (33%)

Net profit ( cash flow - tax)

Net Cash Flow ( Net profit after tax + depreciation)

Discounting factor (7.73%)

Discounting factor (7.73%)

Net Present value = NCF * DF

Initial cost

-1,630,000

-49,388

-15,137

Total cost

-1,694,525

-1,694,525

1

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^1

0.93

$ 334,358.48

2

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^2

0.86

$      310,367.10

3

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^3

0.80

$              288,097.19

4

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^4

0.74

$               267,425.22

5

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^5

0.69

$               248,236.53

6

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^6

0.64

$               230,424.70

7

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^7

0.59

$               213,890.94

8

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^8

0.55

$               198,543.52

9

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^9

0.51

$               184,297.34

10

Cost savings

469,222

138,866

330,356

109,018

221,339

360,204

1/(1+0.0773)^10

0.47

$               171,073.37

NPV

NPV ( sum total)

$               752,189.39

To calculated the cash flow, we need to add back the depreciation ,as it is anon cash expenses.

Net Present Value = Total Cash Inflows from Investments – Cost of Investments

So we can see that NPV= $752,189.32